How do I find the equation of line parallel to the line 3x-4y=-4
Answer by jim_thompson5910(35256)

Solve for y

3x-4y=-4

-4y=-4-3x

-4y=-3x-4

y=(-3x-4)/(-4)

y=(-3x)/(-4)-4/(-4)

y = (3/4)x + 1

The final equation listed above is in the form y = mx+b where m = 3/4

The slope is m = 3/4

Any line with a slope of m = 3/4 will be parallel to the given line as long as the y-intercept of this other line isn't b = 1. For example, y = (3/4)x + 10 is parallel to the given line (slope = 3/4, y-intercept = 10)